lagrangian

a pendulum consists of a mass m suspended by a masslessspring with unextended length b and spring constant k. thependulum's point of support rises verticallywith constantacceleration a.
a. use the lagrangian method to find the equations of motion
b. determine the Hamilton's equation of motion
c. What is the period of small oscillations?
Answer #1

Draw the free diagram of the pendulum. First write the lagrangain equation. L=T-U ....... (1) Here, T is the kinetic energy a

From the figure. tan = And. x=(b+r) sin e y=(b+r)cos e Let b+r=r Therefore, i=rsin e +(6+r) cose() And, y = rcos 8-(+r)sin (0

Determine the potential energy. U=-mgy + kr ? =-mg (b+r) cos 2 + 2 ka? Substitute the value of kinetic energy and potential e

Hence, the required equations of motions are mg cose-lor= mil and (b) Calculate the vertical acceleration. y = (6+r)cos e = [

m 1 j=rcos 8 – rsin e()-(b+r)cos (82)-(6+r) sin e(ö) -- cos - 2r sin (0)-(6+r)cos ( 22)-(6+r)sine -7 (b+r) =g-rcos 8–2rsin e(

Calculate the value of p. = m(b+r)* è=_Po m(b+r)? Therefore, H = p.r+ p. 6-1 = p (%) n (betri Pe-L t he in-mg (b +r)cose+bein

And, P. Or =-1+mg cose Fore. •-OH P amber And, Pe = m(b+r)? 8 = m(b+r)? Ö=-mg(6+r) sin e Ö=_g sin (6+r)
Answer #2

(c) Ans:

as the point of support rises vertically with constant acceleration

effective gravity , geff = g + a

for the time constant

T = 2pi * sqrt(m/k)

the time period will be same

time period is 2pi * sqrt(m/k)